Copier A can make 300 photocopies of a page in 15 minutes.COpier B can do the same job in 12 minutes.If both copiers work together,how long will it take them to do the job?
Found 2 solutions by mananth, Jk22:
Answer by mananth(16946)

Copier A can make 300 photocopies of a page in 15 minutes.COpier B can do the same job in 12 minutes.If both copiers work together,how long will it take them to do the job?
...
copier A 300 in 15 minutes
Copier A 20 copies in 1 minute
...
copier B 300 in 12 minutes
Copier B 25 copies in 1 minute.
..
Together they can do 20 +25 = 45 copies in 1 minute.
..
300 photocopies
300/45= 6.6 minutes
...
m.ananth@hotmail.ca

Answer by Jk22(389)

It will take : 1/(1/12 + 1/15) = 3*4*5/(5+4) = 20/3 = 6 + 2/3 minutes or 6 minutes 40 seconds.

The reasoning can be done like this : the copy-speed of A is 300/15 pages/minutes

B is 300/12 pages/minute, if they work together, the total speed is 300/15+300/12=20+25=45 page minutes, so for the 300 pages, it will take : 300/45 = 6,666 minutes, or 6 2/3 minutes.